Charles Lackey operates a bakery in Idaho​ Falls, Idaho. Because of its excellent product and excellent​ location, demand has increased by 25​% in the last year. On far too many​ occasions, customers have not been able to purchase the bread of their choice. Because of the size of the​ store, no new ovens can be added. At a staff​ meeting, one employee suggested ways to load the ovens differently so that more loaves of bread can be baked at one time. This new process will require that the ovens be loaded by​ hand, requiring additional manpower. This is the only production change that will be made in order to meet the increased demand. The bakery currently makes 1500 loaves per month. Employees are paid ​$8 per hour. In addition to the labor​ cost, Charles also has a constant utility cost per month of ​$850 and a per loaf ingredient cost of ​$0.35.

Current multifactor productivity for 640 work hours per month​ = ____ ​loaves/dollar ​(round your response to three decimal​ places).

After increasing the number of work hours to 800 per month​, the multifactor productivity​ = ____ ​loaves/dollar ​(round your response to three decimal​ places).

The percentage increase in productivity​ = ____% ​(enter your response as a percentage rounded to two decimal​ places)
Answer #1

Current multi-factor productivity = 1500/(640*8 + .35*1500 + 850)

Current multi-factor productivity =  .231 loaves / dollar

---

New output = 1500*1.25 = 1875

New multi-factor productivity = 1875/(800*8 + .35*1875 + 850)

New multi-factor productivity = .237 loaves / dollar

---

% increase in productivity = (.237-.231)/.231

% increase in productivity = 2.6%